# On automorphism groups of smooth hypersurfaces

### v1

@@ 5. Refined bounds via special monomials
In this section, based on canonical bounds, we derive several refined upper bounds for the size of automorphism groups of smooth forms containing monomials involving different primitive constituents (Lemmas 5.4, 5.5, 5.6, 5.7). Such bounds will be frequently used in Section 7.
@@
Recall that $F^{(d-1,1)}\in S^{d-1}(V_{1}^{*})\otimes S^{1}(V_{2}^{*})$. Let $W_{2}^{\prime}$ be the minimal subvector space of $V_{2}^{*}$ such that $F^{(d-1,1)}\in S^{d-1}(V_{1}^{*})\otimes S^{1}(W_{2}^{\prime})$. By $F^{(d-1,1)}\neq 0$, we have $W_{2}^{\prime}\neq 0$. Since $G$ preserves $F$ and $V_{i}$ $(i=1,2)$ are $G$-stable, we have $G$ preserves $F^{(d-1,1)}$, which implies $W_{2}^{\prime}$ is $G$-stable. From the irreducibility of $V_{2}$, we get that $V_{2}^{*}$ is also irreducible under the (right) action by $G$. Then $W_{2}^{\prime}=V_{2}^{*}$, which implies a decomposition of $F^{(d-1,1)}$ as in the lemma. ∎
For our purposes and conventions, we recall some notations in $\bf{Set\textrm{-}up}$ 3.1.
@@
Let $F=F(x_{1},\dots,x_{r})$ be a smooth form of degree $d\geq 3$ with $r\geq 3$. Let $G$ be a subgroup of ${\rm Aut}(F)$. Let $V_{i},W_{ij},K_{i},k_{i},r_{ij},H_{ij},m,P,\widetilde{H}_{ij},\phi_{ij},\pi_{ij},\phi,N$ be as in $\bf{Set\textrm{-}up}$ 3.1 and Definition 3.2. Let $\tilde{\bar{H}}_{ij}:=\phi_{ij}(P)$ and $\bar{H}_{ij}:=\pi_{ij}(\tilde{\bar{H}}_{ij})$. We assume that the underlying basis of ${\mathbb{C}}^{r}$ is compatible with the decomposition ${\mathbb{C}}^{r}=\oplus_{i,j}W_{ij}$. The decomposition
@@
Note that $\tilde{\bar{H}}_{ij}$ and $\bar{H}_{ij}$ are normal subgroups of $\widetilde{H}_{ij}$ and $H_{ij}$ respectively since $P$ is normal in $G$. Clearly primitive constituents $H_{ij}$ of $G$ preserve $F_{ij}$. In the following lemmas and their proofs, we adopt the notations in $\bf{Set\textrm{-}up}$ 5.3.
@@
We have the following commutative diagram
@@
$B={\rm diag}(B_{11},\dots,B_{1k_{1}},\dots,B_{m1},\dots,B_{mk_{m}})\in H,$
we have the commutator $C=[A,B]$ is the identity matrix, by the shape of $A$, $B$, $C(F)=F$ and Lemma 5.2.
@@
Thus $A_{2j}\in\langle\{\lambda I_{r_{2j}}|\lambda\in{\mathbb{C}}^{\times}\}\rangle$ for any $1\leq j\leq k_{2}$. Therefore, $A\in N$ and the lemma follows. ∎
By similar arguments in Lemma 5.4, we obtain the following lemma.
@@
$|G|\leq\frac{B(G,F)}{[H_{i_{1}j_{1}}:\bar{\bar{H}}_{i_{1}j_{1}}]}\leq\frac{R(l(G),d)|\bar{\bar{H}}_{i_{1}j_{1}}|\cdot d^{r}\cdot r!}{{\rm JC}(r_{i_{1}j_{1}})}.$
@@
where the last morphism is given by $g\mapsto([g|W_{12}],\ldots,[g|W_{1k_{1}}],[g|W_{21}],\ldots,[g|W_{mk_{m}}])$ and $\widetilde{N}_{11}$ is defined as its kernel. Let $\bar{\bar{H}}_{11}:=\phi_{11}(\widetilde{N}_{11})$. Then $\bar{\bar{H}}_{11}\vartriangleleft\bar{H}_{11}$ and we have the exact sequence
@@
    If $c=1$, $r_{i_{2}j_{2}}=1$ for $(i_{2},j_{2})\in\mathcal{A}$ and $G$ contains $\mu_{r,d}$, then $F_{i_{1}j_{1}}$ is not zero and $H_{i_{1}j_{1}}$ admits an $F_{i_{1}j_{1}}$-lifting preserving some non-zero form of degree $d-1$.
@@
For (2), let $A={\rm diag}(\lambda_{11}I_{r_{11}},\ldots,\lambda_{1k_{1}}I_{r_{1k_{1}}},\ldots,\lambda_{m1}I_{r_{m1}},\ldots,\lambda_{mk_{m}}I_{r_{mk_{m}}})\in N$. By $A(F)=F$, we get $\lambda_{11}^{d-1}\lambda_{i_{2}j_{2}}=\cdots=\lambda_{11}^{d-1}\lambda_{i_{c+1}j_{c+1}}=1$, where $\mathcal{A}=\{(i_{2},j_{2}),\ldots,(i_{c+1},j_{c+1})\}$. Thus, $\lambda_{i_{2}j_{2}}=\cdots=\lambda_{i_{c+1}j_{c+1}}$, which implies $|N|\leq d^{s-(c-1)}$ by Theorem 3.4. From this, we obtain the inequalities in (2) using the associated exact sequences (3.1) and (3.2).
@@
which contradicts the smoothness of $F$ by [OY19, Lemma 3.2]. Therefore, $F_{11}\neq 0$. Since the $S^{d-1}(W_{11}^{*})\otimes S^{1}(W_{i_{2}j_{2}}^{*})$-component $F_{11i_{2}j_{2}}$ of $F$ is not zero, we write $F_{11i_{2}j_{2}}=F_{11}^{\prime}\cdot L_{i_{2}j_{2}}^{\prime}$, where $F_{11}^{\prime}\in S^{d-1}(W_{11}^{*})$ and $L_{i_{2}j_{2}}^{\prime}\in S^{1}(W_{i_{2}j_{2}}^{*})$. By $\mathcal{A}=\{(i_{2},j_{2})\}$, we have $W_{11}\oplus W_{r_{r_{2}j_{2}}}$ is ${\rm Stab}_{G}(W_{11})$-stable. Let $M:=\{(g_{1},g_{2})|\;g_{1}=g|W_{11},g_{2}=g|W_{i_{2}j_{2}},g\in{\rm Stab}_{G}(W_{11})\}$. If $A=(\lambda_{1}I_{r_{11}},\lambda_{2})\in M$, then by $A(F)=F$, we have $\lambda_{1}^{d}=\lambda_{1}^{d-1}\lambda_{2}=1$ and $\lambda_{1}=\lambda_{2}$. Consider the surjective homomorphisms $\varphi_{1}:M\longrightarrow H_{11}$, $(g_{1},g_{2})\mapsto[g_{1}]$ and $\varphi_{2}:M\longrightarrow\widetilde{H}_{i_{2}j_{2}}\cong C_{d}$, $(g_{1},g_{2})\mapsto g_{2}$. For any $(g_{1},1)\in{\rm Ker}(\varphi_{2})$, we have $g_{1}$ preserves both $F_{11}$ and $F_{11i_{2}j_{2}}$, which implies that $\varphi_{1}|{\rm Ker}(\varphi_{2}):{\rm Ker}(\varphi_{2})\rightarrow H_{11}$ is injective. Clearly $|H_{11}|=|M|/d=|{\rm Ker}(\varphi_{2})|$. Thus $\{g_{1}|\;(g_{1},1)\in{\rm Ker}(\varphi_{2})\}\cong H_{11}$ is a desired $F_{11}$-lifting of $H_{11}$. This completes the proof of the lemma. ∎
@@ 6. Primitive groups of automorphisms
Besides the values of ${\rm JC}(r)$, our proof of Theorem 6.1 relies on explicit list of the finite primitive subgroups in ${\rm PGL}(r,{\mathbb{C}})$ for $r\in\{2,3,4,5,6,8\}$, see Blichfeldt [Bli17] for $r\leq 4$ (see also [CS19, Appendix A] for diagrams indicating inclusions between groups in the case $r=4$); Brauer [Bra67] for $r=5$; Lindsey [Lin71] for $r=6$; Feit [Fei76] for $r=8$. For ease of reference, we present a list of finite primitive groups $H$ in ${\rm PGL}(r,{\mathbb{C}})$ with $|H|\geq 3^{r-1}r!$ and $r\in\{2,3,4,5,6\}$ (see Table 1).
@@
Let $\widetilde{H}\subset{\rm GL}(r,{\mathbb{C}})$ be a finite primitive group. Let $H:=\pi(\widetilde{H})$. If $\widetilde{H}\cong H$, then the following statements hold:
@@
By a theorem of Schur, for a finite group $G$, there is a finite group $SC(G)$ such that $SC(G)$ is a central extension of the Schur multiplier $H^{2}(G,{\mathbb{C}}^{\times})$ by $G$ and every projective representation of $G$ can be lifted to $SC(G)$ (cf. [Rot95, Chapter 7]). We call $SC(G)$ a Schur cover group of $G$. If $G$ is perfect, then $SC(G)$ is perfect and unique up to isomorphism (cf. [Rot95, Corollary 11.12]).
@@
For $(n,d)$ equal to $(1,4)$, $(1,5)$, $(1,6)$, and $(1,7)$, the results are already known, see [Pam13] and [Har19]; for $(n,d)=(1,6)$ and $|{\rm Lin}(X_{F})|=216$, see also [BB22]; for $(n,d)=(1,3)$, see e.g. [YYZ23, Lemma 3.11]. For $2\leq n\leq 4$, Fermat cubic $n$-fold has the largest possible order for the automorphism group among all smooth cubic $n$-folds (see [Seg42], [Hos97], [Dol12] for $n=2$; [WY20] for $n=3$; [LZ22], [YYZ23] for $n=4$). For the case $(n,d)=(2,4)$, $F$ with primitive ${\rm Lin}(X_{F})$ of order $\geq 4^{3}\cdot 4!=1536$ is unique (up to isomorphism) and $|{\rm Lin}(X_{F})|=1920$ (see [Bur55, Exercise 6, Chap. XVII], [AOT24, Theorem A]).
@@
For $(n,d)=(2,6)$, we may assume $|H|\geq 6^{3}\cdot 4!=5184$. We then need to consider $H$ to be one of the following groups: $A_{5}^{2}\rtimes C_{2}$, $C_{2}^{4}\rtimes A_{6}$, $C_{2}^{4}.S_{6}$, ${\rm PSP}(4,3).$ Suppose $H\cong A_{5}^{2}\rtimes C_{2}$. Then $H$ has a perfect subgroup $H_{1}\cong A_{5}^{2}$. The Schur cover group of $A_{5}$ (resp. $A_{5}^{2}$) is ${\rm SL}(2,5)$ (resp. ${\rm SL}(2,5)^{2}$) of order $120$ (resp. $14400$). From the character table of ${\rm SL}(2,5)$ (see [CCNPW]), we infer that for each irreducible representation $\rho:{\rm SL}(2,5)^{2}\rightarrow{\rm GL}(4,{\mathbb{C}})$, there is $A\in\rho({\rm SL}(2,5)^{2})$ with $A$ of order $5$ and ${\rm tr}(A)=-1$, which contradicts Lemma 6.3. Thus, $H\ncong A_{5}^{2}\rtimes C_{2}$. Similarly, we have $H\ncong C_{2}^{4}\rtimes A_{6}$, $C_{2}^{4}.S_{6}$, ${\rm PSP}(4,3)$ by choosing $H_{1}$ as $A_{6}$, $A_{6}$, ${\rm PSP}(4,3)$ respectively.
For $(n,d)=(2,8),(2,10)$, we may assume $|H|\geq 8^{3}\cdot 4!=12288$. Thus we only need to consider $H\cong{\rm PSP}(4,3)$, which preserves no forms of degree $d=8,10$ by the Molien’s series of its Schur cover (cf. [ST54, Page 287]), a contradiction by Lemma 6.3.
For $(n,d)=(4,4)$, since $|H|\geq 4^{5}\cdot 6!=737280$, we have $H\cong{\rm PSU}(4,3)$ or $H\cong{\rm PSU}(4,3).C_{2}$. As in the previous cases, these two groups can be ruled out by the Molien’s series of the Schur cover of ${\rm PSU}(4,3)$ (cf. [Tod50, Page 86]) and the fact ${\rm PSU}(4,3)<{\rm PSU}(4,3).C_{2}$.
For $(n,d)=(4,6)$, there is a unique form $F=F(x_{1},\ldots,x_{6})$ of degree $6$ with primitive $\pi({\rm Aut}(F))\cong{\rm PSU}(4,3).C_{2}$ of order $6531840$ ([Tod50]). Smoothness of $F$ can be verified by Magma ([BCP97]). ∎

### v2

@@ 5. Refined bounds via special monomials
In this section, based on canonical bounds, we derive several refined upper bounds for the size of automorphism groups of smooth forms containing monomials involving different primitive constituents (Lemmas 5.4, 5.5, 5.6, 5.7). The motivation for such bounds is to control the size of the automorphism group ${\rm Aut}(F)$ of a smooth form $F$ when ${\rm Aut}(F)$ is not primitive and they will be frequently used in Section 7.
@@
Recall that $F^{(d-1,1)}\in S^{d-1}(V_{1}^{*})\otimes S^{1}(V_{2}^{*})$. Let $W_{2}^{\prime}$ be the minimal subvector space of $V_{2}^{*}$ such that $F^{(d-1,1)}\in S^{d-1}(V_{1}^{*})\otimes S^{1}(W_{2}^{\prime})$. Since $F^{(d-1,1)}\neq 0$, we have $W_{2}^{\prime}\neq 0$. Since $G$ preserves $F$ and $V_{i}$ $(i=1,2)$ are $G$-stable, we have $G$ preserves $F^{(d-1,1)}$, which implies $W_{2}^{\prime}$ is $G$-stable. From the irreducibility of $V_{2}$, we get that $V_{2}^{*}$ is also irreducible under the (right) action by $G$. Then $W_{2}^{\prime}=V_{2}^{*}$, which implies a decomposition of $F^{(d-1,1)}$ as in the lemma. ∎
For our purposes and conventions, we recall some notations in $\bf{Set\textrm{-}up}$ 3.2.
@@
Let $F=F(x_{1},\dots,x_{r})$ be a smooth form of degree $d\geq 3$ with $r\geq 3$. Let $G$ be a subgroup of ${\rm Aut}(F)$. Let $V_{i},W_{ij},K_{i},k_{i},r_{ij},H_{ij},m,s,P,\widetilde{H}_{ij},\phi_{ij},\pi_{ij},\phi,N$ be as in $\bf{Set\textrm{-}up}$ 3.2 and Definition 3.3. Let $\tilde{\bar{H}}_{ij}:=\phi_{ij}(P)$ and $\bar{H}_{ij}:=\pi_{ij}(\tilde{\bar{H}}_{ij})$. We assume that the underlying basis of ${\mathbb{C}}^{r}$ is compatible with the decomposition ${\mathbb{C}}^{r}=\oplus_{i,j}W_{ij}$. The decomposition
@@
Note that $\tilde{\bar{H}}_{ij}$ and $\bar{H}_{ij}$ are normal subgroups of $\widetilde{H}_{ij}$ and $H_{ij}$ respectively, since $\phi_{ij}$ and $\pi_{ij}$ are surjective and $P$ is normal in $G$. Clearly primitive constituents $H_{ij}$ of $G$ preserve $F_{ij}$. In the following lemmas and their proofs, we adopt the notations in $\bf{Set\textrm{-}up}$ 5.3. From the associated exact sequences (3.1) and (3.2), we have
$|G|\leq\frac{B(G,F)}{\prod_{i}[S_{k_{i}}:K_{i}]}\cdot\frac{|N|}{d^{s}}\cdot\frac{|P/N|}{\prod_{i,j}|H_{ij}|}.$ (5.1)
@@
For all $1\leq i\leq m$ and for all $1\leq j\leq k_{i}$, we have the following commutative diagram:
@@
$B={\rm diag}(B_{11},\dots,B_{1k_{1}},\dots,B_{m1},\dots,B_{mk_{m}})\in P,$
[…]
@@
Thus $A_{2j}\in\langle\{\lambda I_{r_{2j}}|\lambda\in{\mathbb{C}}^{\times}\}\rangle$ for any $1\leq j\leq k_{2}$. Therefore, $A\in N$ and $N=\widetilde{N}$. Then $|P/N|=|P/\widetilde{N}|\leq\frac{\prod_{i,j}|H_{ij}|}{\prod_{1\leq j\leq k_{2}}|H_{2j}|}$ and the lemma follows from (5.1). ∎
By similar arguments in Lemma 5.4, we obtain the following lemma using (5.1).
@@
Note that the spaces $\oplus_{(i,j)\neq(a,2)}W_{ij}$ and $W_{a2}$ are ${\rm Stab}_{G}(W_{a2})$-stable subspaces of ${\mathbb{C}}^{r}$. Since $W_{a2}$ is an irreducible ${\rm Stab}_{G}(W_{a2})$-stable subspace, we can apply Lemma 5.2 as in the proof of Lemma 5.4 to prove that $N$ is equal to the kernel of the morphism $P\longrightarrow\prod_{(i,j)\neq(a,2)}H_{ij}$, $g\mapsto([g|W_{ij}])$. From this, we have $|P/N|\leq\frac{\prod_{i,j}|H_{ij}|}{|H_{a2}|}$ in Lemma 5.5 (1) and (2). For $k_{a}\geq 3$, if $[S_{k_{a}}:K_{a}]\geq 2$, then (2) follows; if $[S_{k_{a}}:K_{a}]=1$, then $\frac{|N|}{d^{s}}\leq\frac{1}{d^{k_{a}-1}}$ (similar to the proof of Lemma 5.7 (2)) and we are done by (5.1).
@@
$|G|\leq\frac{B(G,F)}{[H_{i_{1}j_{1}}:\bar{\bar{H}}_{i_{1}j_{1}}]}\leq\frac{R(l(G),d)\cdot|\bar{\bar{H}}_{i_{1}j_{1}}|\cdot d^{r}\cdot r!}{{\rm JC}(r_{i_{1}j_{1}})}.$
@@
where the last morphism is given by $g\mapsto([g|W_{12}],\ldots,[g|W_{1k_{1}}],[g|W_{21}],\ldots,[g|W_{mk_{m}}])$ and $\widetilde{N}_{11}$ is defined as its kernel. Let $\bar{\bar{H}}_{11}:=\pi_{11}\circ\phi_{11}(\widetilde{N}_{11})$. Then $\bar{\bar{H}}_{11}\vartriangleleft\bar{H}_{11}$ and we have the exact sequence
@@
    If $c=1$, $r_{i_{2}j_{2}}=1$ for $(i_{2},j_{2})\in\mathcal{A}$ and $G$ contains $\mathcal{I}_{r,d}$, then $F_{i_{1}j_{1}}$ is not zero and $H_{i_{1}j_{1}}$ admits an $F_{i_{1}j_{1}}$-lifting preserving some non-zero form of degree $d-1$.
@@
For (2), let $A={\rm diag}(\lambda_{11}I_{r_{11}},\ldots,\lambda_{1k_{1}}I_{r_{1k_{1}}},\lambda_{21}I_{r_{21}},\ldots,\lambda_{m1}I_{r_{m1}},\ldots,\lambda_{mk_{m}}I_{r_{mk_{m}}})\in N$. By $A(F)=F$, we get $\lambda_{11}^{d-1}\lambda_{i_{2}j_{2}}=\cdots=\lambda_{11}^{d-1}\lambda_{i_{c+1}j_{c+1}}=1$, where $\mathcal{A}=\{(i_{2},j_{2}),\ldots,(i_{c+1},j_{c+1})\}$. Thus, $\lambda_{i_{2}j_{2}}=\cdots=\lambda_{i_{c+1}j_{c+1}}$, which implies $|N|\leq d^{s-(c-1)}$ by Theorem 3.5. From this, we obtain the inequalities in (2) using the associated exact sequences (3.1) and (3.2).
@@
which contradicts the smoothness of $F$ by [OY19, Lemma 3.2]. Therefore, $F_{11}\neq 0$. Since the $S^{d-1}(W_{11}^{*})\otimes S^{1}(W_{i_{2}j_{2}}^{*})$-component $F_{11i_{2}j_{2}}$ of $F$ is not zero, we write $F_{11i_{2}j_{2}}=F_{11}^{\prime}\cdot L_{i_{2}j_{2}}^{\prime}$, where $F_{11}^{\prime}\in S^{d-1}(W_{11}^{*})$ and $L_{i_{2}j_{2}}^{\prime}\in S^{1}(W_{i_{2}j_{2}}^{*})$. By $\mathcal{A}=\{(i_{2},j_{2})\}$, we have $W_{11}\oplus W_{i_{2}j_{2}}$ is ${\rm Stab}_{G}(W_{11})$-stable. Let $M:=\{(g_{1},g_{2})|\;g_{1}=g|W_{11},g_{2}=g|W_{i_{2}j_{2}},g\in{\rm Stab}_{G}(W_{11})\}$. If $A=(\lambda_{1}I_{r_{11}},\lambda_{2})\in M$, then by $A(F)=F$, we have $\lambda_{1}^{d}=\lambda_{1}^{d-1}\lambda_{2}=1$ and $\lambda_{1}=\lambda_{2}$. Consider the surjective homomorphisms $\varphi_{1}:M\longrightarrow H_{11}$, $(g_{1},g_{2})\mapsto[g_{1}]$ and $\varphi_{2}:M\longrightarrow\widetilde{H}_{i_{2}j_{2}}\cong C_{d}$, $(g_{1},g_{2})\mapsto g_{2}$. For any $(g_{1},1)\in{\rm Ker}(\varphi_{2})$, we have $g_{1}$ preserves both $F_{11}$ and $F_{11i_{2}j_{2}}$, which implies that $\varphi_{1}|{\rm Ker}(\varphi_{2}):{\rm Ker}(\varphi_{2})\rightarrow H_{11}$ is injective. Clearly $|H_{11}|=|M|/d=|{\rm Ker}(\varphi_{2})|$. Thus $\{g_{1}|\;(g_{1},1)\in{\rm Ker}(\varphi_{2})\}\cong H_{11}$ is a desired $F_{11}$-lifting of $H_{11}$. This completes the proof of the lemma. ∎
@@ 6. Primitive groups of automorphisms
Besides the values of ${\rm JC}(r)$, our proof of Theorem 6.1 relies on explicit list of the finite primitive subgroups in ${\rm PGL}(r,{\mathbb{C}})$ for $r\in\{2,3,4,5,6,8\}$, see Blichfeldt [Bli17] for $r\leq 4$ (see also [CS19, Appendix A] for diagrams indicating inclusions between groups in the case $r=4$); Brauer [Bra67, §9] for $r=5$; Lindsey [Lin71, §3] for $r=6$; Feit [Fei76, §2] for $r=8$. For ease of reference, we present a list of finite primitive groups $H$ in ${\rm PGL}(r,{\mathbb{C}})$ with $|H|\geq 3^{r-1}r!$ and $r\in\{2,3,4,5,6\}$ (see Table 1).
@@
Let $\widetilde{H}\subset{\rm GL}(r,{\mathbb{C}})$ be a finite primitive subgroup and $H:=\pi(\widetilde{H})$. If $\widetilde{H}\cong H$, then the following statements hold:
@@
By a theorem of Schur, for a finite group $G$, there is a finite group $SC(G)$ such that $SC(G)$ is a central extension of the Schur multiplier $H^{2}(G,{\mathbb{C}}^{\times})$ by $G$ and every projective representation of $G$ can be lifted to $SC(G)$ (see [Rot95, Chapter 7]). We call $SC(G)$ a Schur cover group of $G$. If $G$ is perfect (i.e., $G$ is equal to its commutator subgroup), then $SC(G)$ is perfect and unique up to isomorphism (see [Rot95, Corollary 11.12]).
@@
For $(n,d)$ equal to $(1,4)$, $(1,5)$, $(1,6)$, and $(1,7)$, the results are already known, see [Pam13] and [Har19]; for $(n,d)=(1,6)$ and $|{\rm Lin}(X_{F})|=216$, see also [BB22]; for $(n,d)=(1,3)$, see e.g. [YYZ23, Lemma 3.12]. For $2\leq n\leq 4$, Fermat cubic $n$-fold has the largest possible order for the automorphism group among all smooth cubic $n$-folds (see [Seg42], [Hos97], [Dol12] for $n=2$; [WY20] for $n=3$; [LZ22], [YYZ23] for $n=4$). For the case $(n,d)=(2,4)$, $F$ with primitive ${\rm Lin}(X_{F})$ of order $\geq 4^{3}\cdot 4!=1536$ is unique (up to isomorphism) and $|{\rm Lin}(X_{F})|=1920$ (see [Bur55, Exercise 6, Chap. XVII], [AOT24, Theorem A and Corollary B]).
@@
For $(n,d)=(2,6)$, we may assume $|H|\geq 6^{3}\cdot 4!=5184$. We then need to consider $H$ to be one of the following groups: $A_{5}^{2}\rtimes C_{2}$, $C_{2}^{4}\rtimes A_{6}$, $C_{2}^{4}.S_{6}$, ${\rm PSP}(4,3).$ Suppose $H\cong A_{5}^{2}\rtimes C_{2}$. Then $H$ has a perfect subgroup $H_{1}\cong A_{5}^{2}$. The Schur cover group of $A_{5}$ (resp. $A_{5}^{2}$) is ${\rm SL}(2,5)$ (resp. ${\rm SL}(2,5)^{2}$) of order $120$ (resp. $14400$). From the character table of ${\rm SL}(2,5)$ (see [CCNPW, Page 2]), we infer that for each irreducible representation $\rho:{\rm SL}(2,5)^{2}\rightarrow{\rm GL}(4,{\mathbb{C}})$, there is $A\in\rho({\rm SL}(2,5)^{2})$ with $A$ of order $5$ and ${\rm tr}(A)=-1$, which contradicts Lemma 6.3. Thus, $H\ncong A_{5}^{2}\rtimes C_{2}$. Similarly, we have $H\ncong C_{2}^{4}\rtimes A_{6}$, $C_{2}^{4}.S_{6}$, ${\rm PSP}(4,3)$ by choosing $H_{1}$ as $A_{6}$, $A_{6}$, ${\rm PSP}(4,3)$ respectively.
For $(n,d)=(2,8),(2,10)$, we may assume $|H|\geq 8^{3}\cdot 4!=12288$. Thus we only need to consider $H\cong{\rm PSP}(4,3)$, which preserves no forms of degree $d=8,10$ by the Molien’s series of its Schur cover (see [ST54, Page 287]), a contradiction by Lemma 6.3.
For $(n,d)=(4,4)$, since $|H|\geq 4^{5}\cdot 6!=737280$, we have $H\cong{\rm PSU}(4,3)$ or $H\cong{\rm PSU}(4,3).C_{2}$. As in the previous cases, these two groups can be ruled out by the Molien’s series of the Schur cover of ${\rm PSU}(4,3)$ (see [Tod50, Page 86]) and the fact ${\rm PSU}(4,3)<{\rm PSU}(4,3).C_{2}$.
For $(n,d)=(4,6)$, there is a unique form $F=F(x_{1},\ldots,x_{6})$ of degree $6$ with primitive $\pi({\rm Aut}(F))\cong{\rm PSU}(4,3).C_{2}$ of order $6531840$ ([Tod50, §3, §4]). Smoothness of $F$ can be verified by Magma ([BCP97]). ∎